# A Survey on 3D Egocentric Human Pose Estimation

### v1

@@ 3.1 Skeletal Based Methods
In this section, we have provided details on the skeletal based egocentric pose estimation methods. A brief overview of skeletal based methods has been highlighted in Table 2.
Rhodin et al. [21] pioneered a marker-less egocentric motion capture system leveraging a lightweight stereo pair of fisheye cameras integrated into a helmet or virtual reality headset. Their technique utilized a generative pose estimation framework for fisheye views, incorporating a ConvNet-based body-part detector. This method proved especially beneficial for virtual reality applications that require unrestricted movement and interaction while displaying a fully motion-captured virtual body. However, real-time execution was not achieved with this approach.
To solve the egocentric 3D pose estimation problem in real-time, Mo2Cap2[22] employed a two-scale location invariant convolutional network to detect 2D joints, accounting for perspective and radial distortions. It then utilized a location-sensitive distance module to estimate absolute camera-to-joint distances. Actual joint positions were recovered by back-projecting 2D detections, considering distance estimations and fisheye lens properties. This approach guaranteed accurate 3D pose estimation and precise 2D overlay of results. But, their method failed dramatically for scenes with severe occlusions.
@@
Several methods [25, 19, 45] have tried to account for this occlusion problem. EgoGlass [25] addressed a solution for low-visible joints. They proposed a solution that utilizes body part information for enhanced pose detection, particularly in occlusion scenarios. The 2D module includes branches for 2D heatmap and body part information prediction. The 3D module uses a pseudo-limb mask concept to address occlusion challenges in real-world images. The 3D module also serves as an autoencoder for joint heatmaps, aiming to achieve robust 3D body pose estimation and encoding uncertainty in 2D predictions from multiple views. [19] also worked with problems like highly occluded region, closely interacting with the scene in egocentric pose estimation. The proposed method introduces an egocentric depth estimation network for predicting scene depth maps behind the human body using a wide-view egocentric fisheye camera, addressing occlusion caused by the human body through a depth-inpainting network. Additionally, a scene-aware pose estimation network was presented for 3D pose regression. [45] utilized a Vector Quantized-Variational AutoEncoder (VQ-VAE) to predict and optimize human pose, addressing the challenge of obscured lower body appearance in egocentric views. By leveraging a codebook learned from extensive human pose datasets, the method significantly improved generalization performance.
Skeletal Methods Year Highlighted features Dataset Evaluation Metrics
Egocap [21] 2016 First marker-less motion capture system; utilized pose estimation framework for fisheye views with a ConvNet-based body-part detector. EgoCap PCK
Jiang et al. [46] 2017 Leveraged dynamic motion signatures and static scene structures to infer the invisible pose efficiently. custom dataset by Kinect V2 sensor MPJPE
Mo2Cap2 [22] 2019 Real-time; disentangled 3D pose estimation, addressed 2D joint detection, camera-to-joint distances, and joint position recovery for accurate results and a precise 2D overlay. Mo2Cap2 MPJPE
EgoGlass [25] 2021 Utilized body part information for low-visible joints and tackling self-occlusion by preserving uncertainty information. EgoGlass MPJPE, P-MPJPE
xr-EgoPose [24] 2019 Encoder-decoder model for VR headset images, addressing resolution differences in upper and lower body poses, with a dual-branch decoder preserving uncertainty information. xR-EgoPose MPJPE
You2Me [47] 2020 Implied interactions with a visible second person, utilizing dyadic interactions and incorporating dynamic first-person motion features, static scene features along with explicit second-person body pose interactions for robust pose inference. You2Me MPJPE.
Zhang et al. [48] 2021 Incorporated an automatic calibration module with self-correction to alleviate image distortions in fisheye cameras, ensuring consistency between 3D predictions and distorted 2D poses for enhanced accuracy. xR-EgoPose MPJPE
Wang et al. [20] 2021 Spatio-temporal optimization framework that combines 2D and 3D keypoints, VAE-based motion priors, and SLAM-based camera pose estimation, achieving accurate and stable global body pose estimation in egocentric videos. Mo2Cap2, AMASS PA-MPJPE, BA-MPJPE and Global MPJPE
Wang et al. [18] 2022 Employed weak supervision from an external view, leveraging a spatio-temporal optimization technique, and utilized synthetic data with domain adaptation for improved egocentric human pose estimation. EgoPW PA-MPJPE, BA-MPJPE
Akada et al. [17] 2022 Enhances 3D pose estimation by integrating a stereo-based 2D joint location estimation module with weight-sharing encoders and a multi-branch autoencoder for uncertainty capture. UnrealEgo MPJPE, PA-MPJPE
Ego+X [49] 2022 dual-camera framework for 3D global pose estimation and social interaction characterization, leveraging visual SLAM and a Pose Refine Module (PRM) for spatial and temporal accuracy and characterizes social interactions based on global 3D human poses. ECHA MPJPE, PA-MPJPE and Bn-MPJPE
Wang et al. [19] 2023 First scene-aware egocentric pose estimation framework utilizing a depth estimation network to handle occluded regions and close interactions. EgoGTA, EgoPW-Scene MPJPE, PA-MPJPE
EgoFish3D [27] 2023 A self-supervised framework for egocentric 3D pose estimation, utilizing real-world data with three key modules: third person view, egocentric, and interactive modules, achieving accurate results without the need for ground truth annotations. ECHP MPJPE, PA-MPJPE and BA-MPJPE
Ego3DPose [50] 2023 A stereo matcher network and perspective embedding heatmap representation, enhancing accuracy through independent learning of stereo correspondences and leveraging 3D perspective information. UnrealEgo MPJPE, PA-MPJPE
Ego-STAN [51] 2023 Tackles fisheye distortion and self-occlusions in egocentric human pose estimation through a domain-guided spatio-temporal transformer, using 2D image representations, feature map tokens, and 3D pose estimation for accurate joint localization and uncertainty management. xr-EgoPose MPJPE
Dhamanaskar et al. [28] 2023 Utilized third-person view information, creating a self-supervised neural network that establishes a shared space for consistent 3D body pose detection across diverse video settings, ensuring adaptability to real-world scenarios with unknown camera configurations. First2Third-Pose MPJPE
EgoFormer [52] 2023 Leveraged video context and establishing long-term temporal relationships. It addresses ambiguity in first-person videos, surpassing dynamic features, and introduces a novel motion clue representation for enhanced accuracy. CMU Mocap [53] MPJPE, MPJAE.
xR-EgoPose [24], an innovative encoder-decoder architecture was introduced to enhance accuracy considering the differences in resolution between upper body and lower body. This model addresses resolution differences in upper and lower body poses from monocular images captured by VR headset cameras. The dual-branch decoder handled uncertainties in 2D joint locations by first generating 2D heatmaps and then using an autoencoder to regress 3D poses. However, this method struggles when hands fall outside the camera’s range.
@@
To solve this problem of out-of-field-view, [46] developed a method aiming to infer the invisible pose of a person in egocentric videos using dynamic motion signatures and static scene structures. By combining short-term and longer-term pose dynamics, the method utilizes classifiers to estimate pose probabilities and performs joint inference for a longer sequence. In 2021, they extended the idea by proposing another novel technique [54] utilizing both dynamic motion information from camera SLAM and occasionally visible body parts for robust ego pose estimation. It considers the joint estimation of head and body pose, ensuring geometrical consistency and leveraging global coordinate information from camera SLAM. It solves the problem by optimizing the combination of motion and visible shape features. Moreover, EgoTAP [55] addresses out-of-view limbs and self-occlusion issues in stereo egocentric 3D pose estimation by introducing a Grid ViT Heatmap Encoder and Propagation Network. The Grid ViT efficiently summarizes joint heatmaps, preserving spatial relationships crucial for accurate feature embedding. The Propagation Network utilizes skeletal information to predict 3D poses, improving accuracy for both visible and less visible joints.
Figure 3: Illustration of a skeletal model for egocentric 3D pose estimation framework. [52]. It utilizes egocentric videos from a front-facing camera to estimate 3D human pose. Left: An avatar with a head-mounted camera captures synchronized video while performing actions. Right: Output pose sequence generated by EgoFormer.
@@
Scarcity of real-world dataset from egocentric view was a major problem which implicates failure for different proposed methods. To account for this problem, [18] proposed a novel method focusing on enhancing egocentric human pose estimation with weak supervision from an external view. The method introduces a spatio-temporal optimization technique to generate accurate 3D poses for frames in the EgoPW dataset, utilizing them as pseudo labels for training an egocentric pose estimation network. A novel learning strategy is employed to supervise egocentric features with high-quality features from a pretrained external view pose estimation model. It further leverages a synthetic dataset and adopts domain adaptation to minimize the gap between synthetic and real data. [17] also proposed a solution for egocentric pose estimation in an unconstrained environment. They came up with a novel technique to enhance 3D pose estimation by employing a 2D joint location estimation module for stereo inputs. The module utilizes weight-sharing encoders and a decoder leveraging stereo information to boost performance. The 3D module comprises a multi-branch autoencoder, predicting 2D heatmaps to generate 3D pose and reconstructing heatmaps to capture uncertainty.
@@
Perspective distortion is an important problem to consider in egocentric 3D pose estimation as it can cause issues like scale variation, depth ambiguity and limited field of view. To enhance accuracy by tackling this problem, Ego3DPose [50] introduces a Stereo Matcher network that independently learns stereo correspondences and predicts explicit 3D orientation for each limb, avoiding dependence on full-body information. Additionally, a Perspective Embedding Heatmap representation is introduced, allowing the 2D module to extract and utilize 3D perspective information. [51] also addressed the challenges of fisheye distortion and self-occlusions in egocentric human pose estimation by leveraging a domain-guided spatio-temporal transformer model, Ego-STAN. It utilizes 2D image representations and spatiotemporal attention to mitigate distortions and accurately estimate the location of heavily occluded joints. Introducing a spatial concept called feature map tokens (FMT), it draws complex spatio-temporal information from egocentric videos. The model includes a heatmap reconstruction module using deconvolutions, enabling accurate 3D joint coordinate estimation. [48] employed an automatic calibration module with self-correction to mitigate the impact of image distortions on 3D pose estimation. It estimates intrinsic and distortion camera parameters. Unlike traditional post-processing steps, this module ensures consistency between 3D predictions and distorted 2D poses, effectively alleviating the effects of distortions on 3D pose estimation.
@@
When the predicted poses are in the fisheye camera’s local coordinate system instead of the global world coordinate system, it can cause issues like temporal instability. To solve this issue, [20] aimed to estimate accurate and temporally stable global body poses. It employs a spatio-temporal optimization framework, leveraging 2D and 3D keypoints from CNN detection and VAE-based motion priors learned from a mocap dataset. The method addresses challenges like temporal jitters, tracking failures, and unrealistic motions in egocentric pose estimation, by combining local pose estimation, sequential VAEs for motion priors, and SLAM-based camera pose estimation. The method enhances the accuracy and stability in obtaining globally coherent body poses from video sequences. Moreover, Ego+X [49] also proposed a novel egocentric vision framework with two cameras for 3D global pose estimation and social interaction characterization. They proposed Ego-Glo which solves spatial and temporal errors by estimating 3D canonical human pose using a dual-branch network and visual SLAM. Also, their Ego-Soc module performs egocentric social interaction characterization, including object detection and human-human interaction, based on the global 3D human poses.
Generating 3D ground truth data using motion capture system is a cumbersome task. To alleviate this problem, a simple and effective framework, EgoFish3D [27] was introduced for egocentric pose estimation using real-world data and self-supervised learning technique. The proposed method introduces three modules for egocentric 3D pose estimation: a third person view module generating accurate 3D poses from external camera images, an egocentric module predicting 3D poses from a single fisheye image in a self-supervised manner, and an interactive module estimating rotation differences between third person and egocentric views. The method achieves self-supervised egocentric 3D pose estimation without ground truth annotations, utilizing a real-world dataset (ECHP) with synchronized third person and egocentric images.
Yuan et al. [56] introduced a reinforcement learning-based PD control policy for 3D human pose estimation and forecasting from egocentric videos. It learns directly from diverse motion capture data, forecasting stable future motions using a video-conditioned recurrent control technique. Introducing a fail-safe mechanism, their approach outperforms previous methods in controlled and in-the-wild scenarios.
Model Based Methods Year Highlighted features Dataset Evaluation Metrics
@@
Yuan et al. [57] 2018 Integrates control-based modeling, physics simulation, and imitation learning for ego-pose estimation, enabling domain adaptation by considering underlying physics dynamics. CMU Mocap [53] MPJPE, MPJVE
@@
Dittadi et al. [58] 2021 Variational autoencoders for generating human body poses from limited head and hand pose data, addressing challenges through specialized inference models. AMASS MPJPE, MPJVE
@@
CoolMoves [59] 2021 Achieves real-time, expressive full-body motion synthesis for avatars using limited input cues, dynamically fusing stylized examples from skilled performers, excelling in activities like dancing and fighting. CMU MoCap MPJPE, Pearson’s correlation coefficient
@@
EgoRenderer [60] 2021 Renders full-body neural avatars from egocentric fisheye camera images for texture synthesis, pose construction, and neural image translation, addressing challenges of top-down view and distortions. EgoRenderer SSIM, LPIPS, PSNR
HPS [61] 2021 Integrates wearable sensors, IMUs, and a head-mounted camera to accurately track 3D human poses in pre-scanned environments, achieving drift-free motion estimation through fusion of localization and scene constraints. HPS Bidirectional Chamfer Distance
@@
Avatarposer [62] 2022 First learning-based method predicting full-body poses in world coordinates, leveraging transformer encoder, motion input from head and hands, and decoupling global motion for accurate pose estimation. CMU, AMASS MPJPE, MPJRE, MPJVE
@@
FLAG [63] 2022 Flow-based model for realistic 3D human body pose prediction with uncertainty estimates, enhancing prior work through high-quality pose generation and efficient latent variable sampling for optimization. AMASS MPJPE
EgoEgo [64] 2023 Ego body pose estimation using ego head pose estimation leveraging SLAM, and conditional diffusion for disentangled head and body pose estimation. ARES Head Orientation Error, Head Translation Error, MPJPE.
EgoHMR [65] 2023 Scene-conditioned diffusion approach guided by a physics-based collision score, enabling realistic human-scene interactions and accurate estimations for visible body parts while enhancing diversity. EgoBody MPJPE, Vertex-to-Vertex errors
EgoPoser [66] 2023 Used sparse motion sensor; mitigates overfitting with position-invariant prediction, adaptable to diverse body sizes, robust with hands out of view and reduces motion artifacts. AMASS subset [67] MPJPE, MPJVE
@@
Su et al. [68] 2022 A data framework transforms raw video into 3D pose, enriched by a lightweight Self-Perception Excitation (SPE) module for egocentric self-awareness. Mocap dataset [56] MPJPE, MRLE.
@@
SimpleEgo [69] 2024 Directly predicts joint rotations as matrix Fisher distributions, providing robust uncertainty estimation and realistic deployment prospects. SynthEgo MPJPE, PA-MPJPE
@@
Table 3: Popular model shape generation based egocentric 3D pose estimation models.
@@
Recent research shows that linking first-person and third-person view [70, 71, 72] plays a crucial role for better understanding wearer’s action and poses. [73] developed a novel method to make 3D body pose detectors for first-person videos more reliable. They used visual information from paired third-person videos to create a shared space where different views of the same pose are close together. They trained a special neural network to learn this shared space in a self-supervised manner, teaching it to distinguish if two views show the same 3D skeleton. This approach doesn’t need adaptation to work with new, unpaired videos, making it suitable for real-world situations where camera settings are unknown.
@@
You2Me [47] addressed the challenge of estimating the 3D body pose of the camera wearer by leveraging interactions with a visible second person. The key insight is that dyadic interactions between individuals help to learn temporal models for interlinked poses even when one person is largely out of the field view. The method incorporates dynamic first-person motion features, static first-person scene features and second-person’s body pose interaction features to explicitly account for the body pose of the camera wearer.
EgoFormer [52], a tansformer-based model for ego-pose estimation in AR and VR applications, addresses the ambiguity in first-person videos by leveraging video context and establishing long-term temporal relationships. EgoFormer extracts effective temporal features, surpassing dynamic features, and introduces a novel representation for motion clues. Figure 3 shows the proposed architecture for the EgoFormer method.
3.2 Model Based Methods

### v2

@@ 3.1 Skeletal Based Methods
In this section, we have provided details on the skeletal based egocentric pose estimation methods. Table 2 provides a brief overview of 17 such skeletal based methods.
Rhodin et al. [51] introduced a marker-less egocentric motion capture system using fisheye cameras embedded in a helmet or VR headset. The method employs a generative pose estimation framework with a ConvNet-based body part detector, ideal for VR applications needing natural movement and interaction. However, it was not able to attain real-time performance. To solve which, Mo2Cap2 [68] uses a two-scale location invariant convolutional network to detect 2D joints, accommodating perspective and radial distortions. It uses a location-sensitive distance module for estimating absolute camera-to-joint distances, and then recovers actual joint positions by back-projecting 2D detections. However, it struggles in scenes with severe occlusions.
@@
EgoGlass [77] solves the occlusion problem by leveraging body part information for improved pose detection. The 2D module incorporates branches for heatmap and body part prediction, while the 3D module employs a pseudo-limb mask approach to handle occlusion in real-world images. This module also functions as an autoencoder for joint heatmaps, enhancing 3D body pose estimation and capturing uncertainty in 2D predictions across multiple views. [67] introduces an egocentric depth estimation network for predicting scene depth maps behind the human body using a wide-view egocentric fisheye camera, addressing occlusion caused by the human body through a depth-inpainting network. Additionally, a scene-aware pose estimation network was presented for 3D pose regression. [25] used a Vector Quantized-Variational AutoEncoder (VQ-VAE) to predict and optimize human pose, addressing the challenge of obscured lower body appearance. xR-EgoPose [60] and SelfPose [61] uses an encoder-decoder architecture designed to improve accuracy in capturing upper and lower body poses from monocular images obtained via VR headset cameras. [60] employs a dual-branch decoder to address resolution discrepancies between the upper and lower body. It handles uncertainties in 2D joint locations by initially generating 2D heatmaps and subsequently using an autoencoder for 3D pose regression.
@@
To solve the problem of out-of-field-view, [28] developed a method aiming to infer the invisible pose of a person in egocentric videos using dynamic motion signatures and static scene structures. By combining short-term and longer-term pose dynamics, the method utilizes classifiers to estimate pose probabilities and performs joint inference for a longer sequence. They extended the idea [29] by using both dynamic motion information from camera SLAM and occasionally visible body parts for robust ego pose estimation ensuring geometrical consistency. EgoTAP [32] addresses out-of-view limbs and self-occlusion issues in stereo egocentric 3D pose estimation by introducing a Grid ViT Heatmap Encoder and Propagation Network. The Grid ViT efficiently summarizes joint heatmaps, preserving spatial relationships. The Propagation Network utilizes skeletal information to predict 3D poses, improving accuracy for both visible and less visible joints.
@@
To reduce the scarcity of real-world datasets from egocentric view, [66] proposed the use of weak supervision from an external viewpoint. The approach utilizes spatio-temporal optimization to generate accurate 3D poses for frames in the EgoPW dataset, using them as labels for training an egocentric pose estimation network. It also incorporates a synthetic dataset and employs domain adaptation to bridge the gap between synthetic and real data. [3] proposed a solution for egocentric pose estimation in an unconstrained environment. It uses a 2D joint location estimation module for stereo inputs by utilizing weight-sharing encoders and a decoder leveraging stereo information to boost performance. The 3D module comprises a multi-branch autoencoder, predicting 2D heatmaps to generate 3D pose and reconstructing heatmaps to capture uncertainty.
@@
Perspective distortion can cause issues like scale variation, depth ambiguity and limited field of view. To tackle this problem, Ego3DPose [33] introduces a Stereo Matcher network that independently learns stereo correspondences and predicts explicit 3D orientation for each limb, avoiding dependence on full-body information. Additionally, a Perspective Embedding Heatmap representation is introduced, allowing the 2D module to extract and utilize 3D perspective information. [48] addressed the challenges of fisheye distortion and self-occlusions by leveraging a domain-guided spatio-temporal transformer model, Ego-STAN. It utilizes 2D image representations and spatiotemporal attention to mitigate distortions and accurately estimate the location of heavily occluded joints. [76] employed an automatic calibration module with self-correction to mitigate the impact of image distortions on 3D pose estimation. Unlike traditional post-processing steps, this module ensures consistency between 3D predictions and distorted 2D poses.
@@
When the predicted poses are in the fisheye camera’s local coordinate system instead of the global coordinate system, it can cause issues like temporal instability. To solve this issue, [64] proposed a method for precise and stable global body pose estimation in egocentric videos. It utilizes CNN-detected 2D and 3D keypoints, VAE-based motion priors, and SLAM-based camera pose estimation. This approach effectively tackles challenges like temporal jitters and tracking failures, significantly enhancing accuracy and stability in obtaining coherent body poses. Ego+X [38] proposed a framework with two cameras for 3D global pose estimation and social interaction characterization. The Ego-Glo module solves spatial and temporal errors using a dual-branch network and visual SLAM. Whereas, the Ego-Soc module performs egocentric social interaction characterization, including object detection and human-human interaction, based on the global 3D human poses.
Body Shape based Methods Year Highlighted Characteristics Dataset Limitations Code/Project Website Link
@@
Yuan et al. [71] 2018 Integrates control-based modeling, physics simulation, and imitation learning for ego-pose estimation, enabling domain adaptation by considering underlying physics dynamics. CMU Mocap [1] Indirect 2D evaluation may not capture full 3D accuracy; limited behaviors may hinder complex motion generalization.
@@
Dittadi et al. [14] 2021 Variational autoencoders for generating human body poses from limited head and hand pose data; addressing challenges through specialized inference models. AMASS [44] Incomplete utilization of temporal history, constraints on body shape variation and reliance on assumed availability of hand signals.
@@
CoolMoves [2] 2021 Achieves real-time, expressive full-body motion synthesis for avatars using limited input cues, dynamically fusing stylized examples from skilled performers, excelling in activities like dancing and fighting. CMU MoCap Limited sensing of legs and feet, resulting in lower body reconstruction jitters and reduced accuracy in foot-driven motions.
@@
EgoRenderer [21] 2021 Renders full-body neural avatars from egocentric fisheye images - texture synthesis, pose construction, and neural image translation; addresses challenges of top-down view and distortions. EgoRenderer Incomplete joint estimation; unnatural motions in SMPL model animations and temporal instability in frame predictions. Project
HPS [20] 2021 Integrates wearable sensors, IMUs and a head-mounted camera for precise 3D pose tracking in pre-scanned environments, eliminating drift with localization and scene constraints. HPS Lack of features and scene changes between static 3D scans and real images. Project
@@
Avatarposer [30] 2022 First learning-based method predicting full-body poses in world coordinates, leveraging transformer encoder and motion input from head and hands CMU Mocap, AMASS Sensitivity to inaccuracies and occlusions in hand tracking data. Code
@@
FLAG [5] 2022 Flow-based model for realistic 3D human body pose prediction with uncertainty estimates, enhancing prior work through high-quality pose generation and efficient latent variable sampling for optimization. AMASS Difficulty in generating complex lower-body poses due to sparse training data and lack of temporal data integration. Project
@@
Su et al. [54] 2022 A data framework transforms raw video into 3D pose, enriched by a lightweight Self-Perception Excitation (SPE) module for egocentric self-awareness. Mocap dataset [72] Dependency on MoCap data and synchronized third-person view videos may limit the method’s real-world applicability.
EgoEgo [35] 2023 Ego body pose estimation using ego head pose estimation leveraging SLAM, and conditional diffusion for disentangled head and body pose estimation. ARES Evaluation on synthetic and relatively small real-world datasets. Code
EgoHMR [75] 2023 Scene-conditioned diffusion approach using a physics-based collision score, realistic human-scene interaction, accurate estimation for visible body parts while enhancing diversity. EgoBody Limited temporal context for reconstructing egocentric human motions. Code
EgoPoser [31] 2023 Used sparse motion sensor; mitigates overfitting with position-invariant prediction, adaptable to diverse body sizes, robust with hands out of view and reduces motion artifacts. AMASS Limited evaluation on diverse real-world scenarios.
@@
SimpleEgo [11] 2024 Directly predicts joint rotations as matrix Fisher distributions, providing robust uncertainty estimation and realistic deployment prospects. SynthEgo Accuracy could be limited when large portions of the body are occluded in the image. Project
@@
Table 3: Popular body shape based egocentric 3D pose estimation models.
@@
Generating 3D ground truth data using motion capture system is a cumbersome task. To alleviate this problem, EgoFish3D [39] proposed three modules: a third-person view module generating accurate 3D poses from external camera images, an egocentric module predicting 3D poses from a single fisheye image via self-supervised learning, and an interactive module estimating rotation differences between third-person and egocentric views. This method achieves self-supervised egocentric 3D pose estimation without ground truth annotations, leveraging a real-world dataset (ECHP) with synchronized third-person and egocentric images. Linking first-person and third-person view [37, 53, 69] plays a crucial role for better understanding wearer’s action and poses. [12] used visual information from paired third-person videos to create a shared space where different views of the same pose are close together. They trained a special neural network to learn this shared space in a self-supervised manner, teaching it to distinguish if two views show the same 3D skeleton.
@@
EgoFormer [36], a tansformer-based model for ego-pose estimation in AR and VR applications, addresses the ambiguity in first-person videos by leveraging video context and establishing long-term temporal relationships. It extracts effective temporal features, dynamic features, and introduces a novel representation for motion clues. You2Me [46] addresses the challenge of estimating the 3D body pose of the camera wearer by leveraging interactions with a visible second person. The key insight is that dyadic interactions between individuals help to learn temporal models for interlinked poses even when one person is largely out of the field view. The method incorporates dynamic first-person motion features, static first-person scene features, and second-person’s body pose interaction features to explicitly account for the body pose of the camera wearer.
3.2 Body Shape Based Methods